Solve each problem and provide the complete solution.
note: topic is all about Gravitational force

We are seated on the surface of the earth. What is the gravitational strength of the earth on you if the mass of the earth is 5.9x1024kg and a radius of 6.4x106 m?
1
Expert's answer
2020-11-30T14:56:35-0500

We can find the gravitational strength of the earth from the formula:


"g=\dfrac{GM}{r^2},""g=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot 5.9\cdot10^{24}\ kg}{(6.4\cdot10^6\ m)^2}=9.61\ \dfrac{m}{s^2}."

Answer:

"g=9.61\ \dfrac{m}{s^2}."